# E2ETune: End-to-End Knob Tuning via Fine-tuned Generative Language Model

### v2

@@ 8.4.2. Quality of Labels
Cost Model. The quality of the cost model significantly impacts label quality, as a substantial portion of the training set is derived from it. To provide a comprehensive evaluation of the cost model, we examine two critical aspects: (1) the accuracy of the cost model and (2) the performance of HEBO assisted by the cost model. First, we perform 10-fold cross-validation to train the cost model. For each fold, we use the coefficient of determination ($R^{2}$) to evaluate the cost model’s accuracy on the held-out test set:
(3) $R^{2}=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}},$
where $y_{i}$ is the observed performance metric for the $i$-th sample, $\hat{y}_{i}$ is the predicted performance metric, $\bar{y}$ is the mean of all observed metrics, and $n$ is the total test set observations. Our cost model achieves an average $R^{2}$ score of 0.862 across the 10-fold validation, indicating high reliability. In addition, during label collection, the cost model is trained using observations from all database instances. Therefore, the quality of the configurations gathered by HEBO with the cost model can be observed in the in-schema part of “HEBO + Cost Model” in Figure 3. Compared with the original HEBO, using the cost model only leads to a marginal decrease in database performance, affirming the reliability of the cost model.
@@ 8.5.1. Workload Generation
For OLAP workloads, we employ GPT-4-Turbo to craft approximately 1,100 OLAP queries per database instance. Given the GPT-4-Turbo API’s response time of approximately 20 seconds, and considering our five database instances, the cumulative query generation time amounts to roughly 31 hours, costing approximately \$150. In contrast, generating OLTP workloads incurs no additional expenses.
@@ 8.5.2. Label Collection
Collecting Data via Real Execution. For OLAP workloads, executing the full workload takes between 1 to 20 minutes. Since the original HEBO method requires replaying the workload in each iteration, tuning an OLAP workload across 100 iterations may take 2 to 30 hours. For OLTP workloads, we use a 1-minute stress test for each tuning iteration. As a result, the complete tuning process for an OLTP workload takes approximately 2 hours.
To ensure accurate test results, workloads are tuned sequentially on the server. With each database instance tuning 13 workloads, the full process takes around 30 days on a single machine. In practice, we use eight identical CPU servers to expedite this process.
@@ 9. DISCUSSION
In this section, we first discuss the applicable scenarios for E2ETune. By leveraging the advantages of an end-to-end design, E2ETune significantly enhances tuning efficiency while recommending promising configurations. When encountering a new database engine, it is necessary to re-collect training data and train a new model to tune knobs due to different knobs. Consequently, E2ETune is particularly well-suited for stable or enumerable environments, such as cloud services that maintain uniform hardware configurations and standardized DBMS products. After the initial training, E2ETune proves highly effective in recommending configurations for a diverse range of users utilizing cloud databases. Additionally, to further accelerate data collection and enhance label quality, we can explore and leverage better hyperparameter optimization algorithms, such as HPFSO (Ansótegui et al., 2021), SPBOpt (Sazanovich et al., 2021), and DRE (Khazi et al., 2023), as replacements for HEBO in our data collection framework.
@@ 10. CONCLUSION
This paper explores an innovative method for end-to-end tuning of database knobs. Our emphasis is on utilizing advanced LMs to accurately model the intricate mapping between workload characteristics and promising configurations. To achieve this goal, we introduce a data generation framework designed to automatically produce workloads and configuration labels, serving as training samples for the LM. Through extensive experiments on OLAP and OLTP benchmarks, we demonstrate that E2ETune not only significantly improves the tuning efficiency over existing approaches, but also showcases notable database performance improvements in both in-schema and cross-schema settings.

### v3

@@ 8.4.2. Quality of Labels
Cost Model. The quality of the cost model significantly impacts label quality, as a substantial portion of the training set is derived from it. To provide a comprehensive evaluation of the cost model, we examine two critical aspects: (1) the accuracy of the cost model and (2) the performance of HEBO assisted by the cost model. First, we perform 10-fold cross-validation to train the cost model. For each fold, we use the coefficient of determination ($R^{2}$) to evaluate the cost model’s accuracy on the held-out test set: $R^{2}=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}$, where $y_{i}$ is the observed performance metric for the $i$-th sample, $\hat{y}_{i}$ is the predicted performance metric, $\bar{y}$ is the mean of all observed metrics, and $n$ is the total test set observations. Our cost model achieves an average $R^{2}$ score of 0.862 across the 10-fold validation, indicating high reliability. In addition, during label collection, the cost model is trained using observations from all database instances. Therefore, the quality of the configurations gathered by HEBO with the cost model can be observed in the in-schema part of “HEBO + Cost Model” in Figure 3. Compared with the original HEBO, using the cost model only leads to a marginal decrease in database performance, affirming the reliability of the cost model.
@@ 8.5.1. Workload Generation
For OLAP workloads, we employ GPT-4-Turbo to craft approximately 1,100 OLAP queries per database instance. With a response time of approximately 20 seconds per query and five database instances, the cumulative query generation time amounts to roughly 31 hours, costing approximately \$150. In contrast, generating OLTP workloads incurs no additional expenses.
@@ 8.5.2. Label Collection
Collecting Data via Real Execution. For OLAP workloads, executing the full workload takes between 1 to 20 minutes. Since the original HEBO method requires replaying the workload in each iteration, tuning an OLAP workload across 100 iterations may take 2 to 30 hours. For OLTP workloads, we use a 1-minute stress test for each tuning iteration. As a result, the complete tuning process for an OLTP workload takes approximately 2 hours. To ensure accurate test results, workloads are tuned sequentially on the server. With each database instance tuning 13 workloads, the full process takes around 30 days on a single machine. In practice, we use eight identical CPU servers to expedite this process.
@@ 9. DISCUSSION
In this section, we first discuss the applicable scenarios for E2ETune. By leveraging the advantages of an end-to-end design, E2ETune significantly enhances tuning efficiency while recommending promising configurations. When encountering a new database engine, it is necessary to re-collect training data and train a new model to tune knobs due to different knobs. Consequently, E2ETune is particularly well-suited for stable or enumerable environments, such as cloud services that maintain uniform hardware configurations and standardized DBMS products. After the initial training, E2ETune proves highly effective in recommending configurations for a diverse range of users utilizing cloud databases. We are deploying E2ETune in real-world applications at ByteDance and will open-source our fine-tuned model to support further research. Additionally, to accelerate data collection and enhance label quality, we can explore and leverage better hyperparameter optimization algorithms, such as HPFSO (Ansótegui et al., 2021), SPBOpt (Sazanovich et al., 2021), and DRE (Khazi et al., 2023), as replacements for HEBO in our data collection framework.
@@ 10. CONCLUSION
This paper explores an innovative method for end-to-end database knob tuning. Our emphasis is on utilizing the advanced language model (LM) to accurately capture the complex mapping between the workload and its promising configuration. To achieve this goal, we introduce a data generation framework to automatically produce workloads and their promising configurations, serving as training data samples for the LM. Through extensive experiments on OLAP and OLTP benchmarks, we demonstrate that E2ETune not only significantly improves the tuning efficiency over existing approaches, but also showcases notable database performance improvements in both in-schema and cross-schema settings.
Acknowledgements.
This work is supported by the National Key Research & Development Plan of China (2023YFF0725100) and the National Natural Science Foundation of China (62322214, U23A20299, U24B20144, 62172424, 62276270). We also acknowledge the support of the Public Computing Cloud, Renmin University of China.